# When Ads Become Profiles: Uncovering the Invisible Risk of Web Advertising at Scale with LLMs

### v2

@@ Appendix D Example of Human-Gemini Disagreement
Human–Gemini Disagreement Refer to caption Ad Title   ACT Rebate Finder Ad Text   We Need Homeowners Interested In Trying Solar Energy. Generate your own electricity – you may pay nothing for it. | Eligible participants must: | 1. Own a home… See more Gemini Caption
@@
Descriptive Category Informative, Eco-friendly, Home improvement, Clean energy promotion ✓ Accurate (Yes)   Inaccurate (No) IAB Tier 1 Category Home and Garden Services, Green/Eco, Real Estate ✓ Accurate (Yes)   Inaccurate (No) Key Entities Tesla Powerwall, Solar energy, Home battery system, Homeowners, Electricity generation ✓ Accurate (Yes)   Inaccurate (No)

### v3

@@ Appendix D Example of Human-Gemini Disagreement
Human–Gemini Disagreement Refer to caption Ad Title   ACT Rebate Finder
Ad Text   We Need Homeowners Interested In Trying Solar Energy. Generate your own electricity – you may pay nothing for it. | Eligible participants must: | 1. Own a home… See more
Gemini Caption
@@
Descriptive Category
Informative, Eco-friendly, Home improvement, Clean energy promotion
✓ Accurate (Yes)   Inaccurate (No)
IAB Tier 1 Category
Home and Garden Services, Green/Eco, Real Estate
✓ Accurate (Yes)   Inaccurate (No)
Key Entities
Tesla Powerwall, Solar energy, Home battery system, Homeowners, Electricity generation
✓ Accurate (Yes)   Inaccurate (No)